OpenStudy (abhisindhu):

THE resultant of vector a AND b is perpendicular to the vector a And its magnitude is half that of vector b . find the angle between vector a And b ?

OpenStudy (michele_laino):

please what operation do you mean when you write "a AND b"?

OpenStudy (abhisindhu):

I WROTE a in small so that there wouldn't be any confusion . YOU can take vectors AS P AND Q too

OpenStudy (michele_laino):

please wait, I think that the expression "a AND b" means vector addition, so the problem is solved by these equations: \[\begin{gathered} {\left( {\frac{b}{2}} \right)^2} = {a^2} + {b^2} - 2ab\cos \theta ,\quad \left( {{\text{Theorem of Carnot}}} \right) \hfill \\ \hfill \\ {b^2} = {a^2} + {\left( {\frac{b}{2}} \right)^2},\quad \left( {{\text{Theorem of Pitagora}}} \right) \hfill \\ \end{gathered} \]

OpenStudy (abhisindhu):

CAN you help me by illustrating it in a much better way thanks

OpenStudy (michele_laino):

here is the situation: